A standard die is rolled. Find the probability that the number rolled is less than 33. Express your answer as a fraction in lowest terms or a decimal rounded to the nearest millionth.
Answer #1

Given a standard die is rolled,

S = { 1 ,2 3, 4, 5, 6 }

We have to find the probability that the number rolled is less than 3 is,

number rolled is less than 3 = 1,2

So the probability that the number rolled is less than 3 is, = 2/6 = 1/3 = 0.333333 (nearest millionth)